Find an equation for a line parallel to 4x + 5y + 2 = 0
Found 2 solutions by stanbon, josgarithmetic:
Answer by stanbon(75887)

Find an equation for a line parallel to 4x + 5y + 2 = 0
----------------
1st: Find the slope of the given line.
y = (-4/5)x - (2/5)
slope = -4/5
-------------------
2nd: Slope of the parallel line must be -4/5
------
Ans: y = (-4/5)x + (any number) will be parallel to the given line.
--------------------
Cheers,
Stan H.

Answer by josgarithmetic(39617)

4x+5y+r=0 and r is any real number you want other than 2.